OpenStudy (anonymous):

Solve this system of equations by the substitution method: x+2y=5 -x+y=-5

OpenStudy (anonymous):

add them then y=0, x=5

OpenStudy (anonymous):

x=5-2y -(5-2y)+y -5+2y+y=-5 -5+3y=-5 3y=0 y=0 and x=5